# Scale-DiT: Ultra-High-Resolution Image Generation with Hierarchical Local Attention

### v1

@@ 4 Experiments
Figure 3: 4K comparison with leading baselines. Zoom in to observe the fine details. Scale-DiT produces high-resolution results with superior fidelity. Additional comparisons with other baselines are provided in the Appendix. The middle column shows methods finetuned on 4K data, while the right column shows training-free methods.
@@ 4.1 Implementation Details
Experimental Setup. We adopt FLUX.1-dev Black Forest Labs (2025) as our text-to-image backbone, utilizing the Hugging Face Diffusers library for implementation. Parameter-efficient fine-tuning is conducted using LoRA from the PEFT library, with a rank of 16. Our custom attention mechanisms are implemented following the principles of FlashAttention-2 Dao (2023) and SageAttention Zhang et al. (2025c). The model was trained for 20,000 steps on two NVIDIA A6000 Ada GPUs (48GB VRAM each), using a per-GPU batch size of 1 with gradient accumulation over 4 steps. All subsequent experiments were conducted on this same hardware configuration. For the fine-tuning dataset, we generated 10,000 images at a resolution of $1024\times 1024$ using the base FLUX.1-dev model, prompted by a randomly sampled collection of high-quality text descriptions. The kernel size is designed to be Q-block=128 and K-block=64 (128 for blocksparse Flash Attention) which suits the granularity supported for Zhang et al. (2025c) on NVIDIA A6000 Ada. All experiments were conducted on a cluster of 8 NVIDIA A6000 Ada.
@@
Metrics and Evaluation Protocol. To ensure a comprehensive and diverse evaluation, we generated a benchmark suite of 1,000 high-quality prompts across various categories using GPT-4o. We assess performance using a standard set of metrics: CLIP Score Radford et al. (2021) for prompt-image alignment, and both Fréchet Inception Distance (FID) Heusel et al. (2017) and Inception Score (IS) Salimans et al. (2016) for image quality. The FID score is computed against a reference set of 10,000 real images from the LAION-High-Resolution Schuhmann et al. (2022) dataset. To specifically evaluate the fidelity of local details, we also report patch-based versions of these metrics, $\text{FID}_{\text{patch}}$ and $\text{IS}_{\text{patch}}$, calculated on local image patches. For all comparisons, competing training-free methods were evaluated on the same FLUX.1-dev base model, following their official implementations to ensure fairness.
@@
Table 1: Quantitative comparison at 4K × 4K resolution. The best result is highlighted in bold, while the second-best result is underlined. Scale-DiTconsistently perform competitive results even compared to training-based methods.
@@
SANA 76.31 74.27 16.68 14.02 0.3197
Diffusion-4K 121.85 120.59 14.39 10.77 0.2844
UltraPixel 77.42 70.94 16.98 13.26 0.3251
URAE 67.39 62.56 17.11 12.39 0.3204
FLUX+SR (BSRGAN) 71.39 63.45 17.08 12.87 0.3210
DemoFusion 74.89 66.37 16.23 13.02 0.3187
DiffuseHigh 81.54 73.35 16.15 13.08 0.3175
I-MAX 70.33 65.67 16.50 12.69 0.3211
HiFlow 69.18 63.72 17.13 13.43 0.3113
Scale-DiT (Ours) 67.03 61.78 17.21 13.31 0.3231
@@
4.2 Comparison with State-of-the-Art Methods
We compare our results with methods that can also generate 4K resolution images. These include data-intensive, training-based methods: SANA Xie et al. (2024), Diffusion-4K Zhang et al. (2025b), FLUX+URAE Yu et al. (2025), and UltraPixel Ren et al. (2024); a super-resolution baseline: FLUX+BSRGAN Zhang et al. (2021); and several training-free approaches: DemoFusion Du et al. (2024a), DiffuseHigh Kim et al. (2025), I-MAX Du et al. (2024b), and HiFlow Bu et al. (2025).
Qualitative Comparison. Figure 1 presents qualitative results of Scale-DiT at $4K\times 4K$, $2K\times 3K$, and $2K\times 2K$, demonstrating superior text-to-image generation with high-fidelity details and coherent global structure. A more detailed comparison at 4K is shown in Figure 3, where we include a representative subset of leading baselines for clarity; comprehensive comparisons against the nine methods, as well as additional 2K and 4K results, are provided in the Appendix. At 4K resolution, our method renders anatomically correct hands with exceptional detail as shown in the first row, while SANA produces a blurry image, HiFlow introduces diagonal artifacts, and I-MAX fails to generate the correct number of fingers. In the second and third row, Scale-DiT consistently preserves textural clarity, such as fur and tree structures, surpassing competitors while maintaining global coherence. The fourth row highlights highly detailed eye generation at 4K, where our method clearly captures iris structure, eyelashes, and skin texture with remarkable realism. These examples collectively demonstrate Scale-DiT’s ability to synthesize images that combine fine-grained local details with coherent large-scale composition, validating its performance across diverse high-resolution scenarios.
Quantitative Comparison. Table 1 presents a quantitative comparison of our method against the nine state-of-the-art methods at 4K resolution (please refer to the Appendix for 2K comparisons). The results show that our method achieves highly competitive performance across all metrics, even outperforming training-based methods that rely on extensive 4K-resolution training data. Table 2 demonstrates consistent performance on the GenEval Ghosh et al. (2023) benchmark, confirming that our method’s output quality is resolution-agnostic.
Efficiency and Scalability Analysis. The efficiency gains of our approach are summarized in Table 3 and Figure 4. Focusing on methods built upon the FLUX.1-dev backbone for a fair comparison, our method achieves a $1.5\times$ to $2.7\times$ inference speedup over its counterparts (with all $2.4\times$ speedup per denoising iteration). In terms of memory usage, our method is highly efficient: excluding the base model’s footprint ( 33 GB), it requires only 2.8 GB of additional VRAM, whereas a naive implementation with FlashAttention consumes 14.1 GB, making our approach a lot more memory-efficient, which enables inferencing on smaller consumer-level GPUs, (i.e. 40 GB A100), and more GPU friendly finetuning. Figure 4 further illustrates scalability. As the target resolution increases from 1K to 4K, the gap in both latency and memory usage between our method and the dense attention baseline using FlashAttention2 Dao (2023) widens significantly.
@@ 4.2 Comparison with State-of-the-Art Methods
Compared with Linear Attention, our approach remains faster and more memory-efficient. Linear attention, while computationally faster than FlashAttention, requires scaling the hidden dimension with sequence length to maintain performance Han et al. (2024); Meng et al. (2025); Zhang et al. . In contrast, our method uses local window attention, which does not need to scale with resolution, enabling consistent memory and latency efficiency as resolution increases. Since linear attention requires the hidden dimension to scale with resolution, previous works would need to fine-tune on high-resolution data for each target resolution, whereas our framework achieves ultra-high-resolution synthesis using only 1K-resolution training data without any performance drop. Collectively, these results validate the scalability, efficiency, and robustness of our proposed framework.
Figure 4: Memory and speed statistics when scaling from 1K to 4K comparing with native FLUX.1 Dev with FlashAttention-2, Linear Attention used in SANA, and naive matmul attention. Our method’s advantage in efficiency is more and more significant as resolution increases comparing to the others. OOM stands for out of memory.
Table 3: Comparison of memory (GB) and latency (sec; time to generate one 4K latents (65536 tokens)) across FLUX.1-dev-based 4K generation methods. Our method achieves $2\times$ speedup and significant memory reduction (given that the base model itself is around 33 GB in size).
Method Ours FLUX HiFLOW URAE I-MAX
Memory(GB) 35.8 47.2 43.4 41.3 43.9
Latency(sec) 395 871 634 823 1091
Seconds per iteration(sec/iter) 13.25 31.11 31.59 29.40 35.63
@@ 5 Ablation Study
Window size and resolution ratio. We conduct ablation study on window size (ws) and ratio between high-res and low-res image ($\rho$). We experiment window sizes from 256, 512, and 1024, and $\rho=2,4$. As shown in Table 4, the variations in FiD scores across these configurations were not statistically significant. To maximize computational efficiency without compromising performance, we selected the smallest window size (ws=256) and the largest ratio ($\rho$=4). It is worth noting that smaller window sizes or larger ratios are constrained by the fixed granularity of the kernel which depends on the hardware design of the GPU.
@@ 6 Conclusion
We present Scale-DiT, a novel framework enabling pre-trained diffusion transformers to generate ultra-high-resolution images without requiring additional high-resolution training data. Scale-DiT introduces a hierarchical attention mechanism that partitions high-resolution latents into fixed-size local windows while maintaining global coherence through low-resolution guidance with scaled positional anchors. The framework comprises three key components: (i) efficient local window attention that reduces quadratic complexity to near-linear scaling; (ii) global semantic preservation via low-resolution guidance latents with position scaling; and (iii) parameter-efficient joint denoising through LoRA-based adaptations trained solely on commodity resolutions. Extensive experiments demonstrate that Scale-DiT achieves superior visual quality at 4K resolution while delivering over $2\times$ speedup and less memory usage compared to dense attention baselines, establishing a practical paradigm for scalable ultra-high-resolution text-to-image generation.
Appendix
Appendix A 4K Gallery
In Figure 678 we present more 4K resolution results. More results are in the supplementary materials.
Refer to captionRefer to caption
Figure 6: More 4K results
Refer to captionRefer to caption
@@
Appendix B 2K Gallery
In Figure 9, we present more 2K resolution results. More results are in the supplementary materials.
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
@@
Appendix C LLM Usage Declarations
We declare that Large Language Models (LLMs) were used in limited capacity during the preparation of this manuscript. Specifically, LLMs were employed for: (i) generating diverse text prompts for model evaluation and dataset creation, (ii) grammar checking and language refinement of the manuscript, and (iii) assisting in data collection procedures for experimental validation. All core technical contributions, experimental design, analysis, and conclusions presented in this work are entirely ours own. The use of LLMs did not influence the scientific methodology, results interpretation, or theoretical contributions of this research.
Appendix D REPRODUCIBILITY STATEMENT
We will release our source code upon acceptance.
Appendix E Complete Qualitative Comparison
In figure 10 and figure 11, we present the complete qualitative comparison of 4K resolution images with the quantitatively compared baselines in Table 1. 2K resolution comparison are demonstrated in Figure 12
Figure 10: Complete comparison against the 9 baselines on 4K resolution. Zoom in to view the details, our method produces the best results.
Figure 11: Complete comparison against the 9 baselines on 4K resolution. Zoom in to view the details, our method produces the best results.
Figure 12: Complete comparison against the 9 baselines on 2K resolution. Zoom in to view the details, our method produces the best results.
Appendix F Quantitative Comparison on 2K Resolution
Table 5 demonstrates the quantitative results for $2K\times 2K$ resolution image generation. Our method is in leading position across all metrics.
Table 5: Quantitative comparison at 2K × 2K resolution. The best result is highlighted in bold, while the second-best result is underlined.
[…]
SANA 75.79 74.26 20.15 15.21 0.3188
Diffusion-4K 103.22 97.10 19.88 13.79 0.3159
UltraPixel 75.51 72.97 21.31 16.12 0.3148
URAE 66.24 62.07 21.53 17.32 0.3214
FLUX+SR (BSRGAN) 74.38 72.84 21.07 16.89 0.3201
DemoFusion 78.69 72.45 20.93 15.14 0.3197
DiffuseHigh 71.69 65.58 20.57 15.32 0.3132
I-MAX 67.39 62.78 21.43 17.09 0.3204
HiFlow 67.84 62.06 21.68 16.94 0.3235
Scale-DiT (Ours) 67.23 61.46 21.72 17.26 0.3216
Appendix G Additional implementation details
To enhance the Rectified Flow model’s stability when generalizing to extrapolated resolutions, we adopt inference techniques suggested by previous works Bu et al. (2025); Du et al. (2024b) in the high-resolution generation process, including NTK-aware scaled RoPE.
Appendix H Discussion on Advantages of Local Attention
H.1 Linear Computational Scalability
Our local window attention achieves linear scaling with respect to image area, fundamentally different from the quadratic complexity of full attention. For an image of size $H\times W$ with local windows of size $w\times w$:
$\text{Computational Complexity}=\frac{H\times W}{s^{2}\times w^{2}}\times w^{4}=O(HW)$ (1)
where $s$ is the VAE downsampling factor. This linear relationship means that doubling the image resolution doubles the computational cost, rather than quadrupling it as in full attention systems. The computational density (operations per pixel) remains constant:
$\rho=\frac{\text{Total Operations}}{H\times W}=\frac{w^{2}}{s^{2}}=\text{constant}$ (2)
This property enables practical scaling to arbitrarily large resolutions with predictable resource requirements.
H.2 Quality Preservation Across Resolutions
A critical advantage of our approach is that representational quality remains constant regardless of output resolution. Since every local window operates on the same relative position range $\mathcal{R}_{\text{local}}=\{(\Delta i,\Delta j):|\Delta i|,|\Delta j|\leq w-1\}$, and this range falls entirely within the pretrained distribution $\mathcal{D}_{\text{train}}$, each window achieves identical quality.
The per-unit quality across the entire image is:

### v2

@@ 4.1 Implementation Details
Experimental Setup. We adopt FLUX.1-dev [2] as our text-to-image backbone, utilizing the Hugging Face Diffusers library for implementation. Parameter-efficient fine-tuning is conducted using LoRA from the PEFT library, with both rank and alpha equal to 16. LoRA layers are only applied to the Query, Key, and Value projection layers across the DiT blocks. We adopt pyriology scheduler with default learning rate of 1 and weight decay of 0.01, and set bias correction adn safe-guard warmup to true. During inference, we set the NTK factor of RoPE embedding to be 10 and disabled the dynamic shifting of the scheduler to fir the ultra-high resolution generation following [3, 13]. Our custom attention mechanisms are implemented following the principles of FlashAttention-2 [8] and SageAttention [61]. The model was trained for 20,000 steps on two NVIDIA RTX A6000 Ada GPUs (48GB VRAM each), using a per-GPU batch size of 1 with gradient accumulation over 4 steps. All subsequent experiments were conducted on this same hardware configuration. For the fine-tuning dataset, we generated 10,000 images at a resolution of $1024\times 1024$ using the base FLUX.1-dev model, prompted by a randomly sampled collection of high-quality text descriptions. The kernel size is designed to be Q-block=128 and K-block=64 (128 for blocksparse Flash Attention) which suits the granularity supported for [61] on NVIDIA A6000 Ada. All experiments were conducted on one NVIDIA RTX Pro 6000 Blackwell GPU (96GB VRAM).
Figure 3: 4K comparison with leading baselines based on FLUX.1 (except for SANA). Zoom in to observe the fine details. UltraGen produces high-resolution results with superior fidelity. Additional comparisons with other baselines are provided in the Appendix. The second column section shows methods finetuned on 4K data, the third column section shows training-free method, the right column section shows an other efficient attention method. Our method achieves sharper results than other methods. Notably, although HiFlow produces good results but it shows diagonal artifacts on the first, third and fourth row. We also provide additional results for our method based on FLUX.2, which yields more natural textures.
@@
Metrics and Evaluation Protocol. To ensure a comprehensive and diverse evaluation, we generated a benchmark suite of 1,000 high-quality prompts across various categories using GPT-4o. We assess performance using a standard set of metrics: CLIP Score [37] for prompt-image alignment, CLIP-IQA [47], Fréchet Inception Distance (FID) [18], and Inception Score (IS) [41] for image quality. The FID score is computed against a reference set of 10,000 real images from the LAION-High-Resolution [42] dataset. To specifically evaluate the fidelity of local details, we also report patch-based versions of these metrics, $\text{FID}_{\text{patch}}$ and $\text{IS}_{\text{patch}}$, calculated on local image patches. For all comparisons, competing training-free methods were evaluated on the same FLUX.1-dev base model, following their official implementations to ensure fairness.
4.2 Comparison with State-of-the-Art Methods
We compare our results with methods that can also generate 4K resolution images. These include data-intensive, training-based methods: SANA [50], Diffusion-4K [60], FLUX+URAE [57], UltraFLUX [56], and UltraPixel [38]; a super-resolution baseline: FLUX+BSRGAN [64]; and several training-free approaches: DemoFusion [12], DiffuseHigh [22], I-MAX [13], and HiFlow [3] and an efficient attention method CLEAR [28] (with the 4K ability comes from SDEdit [32] and the optimal setting of window size 32).
Figure 4: Comparison of latency (sec; time to generate one 4K latents (65536 tokens)), time per iteration (sec), and memory (GB) and across FLUX.1-dev-based 4K generation methods. Our method achieves significant speedup and memory reduction.
Qualitative Comparison. Figure 1 presents qualitative results of UltraGen at various resolutions from $2K$ to >8K>8K, demonstrating superior text-to-image generation with high-fidelity details and coherent global structure. A more detailed comparison at 4K is shown in Figure 3, where we include a representative subset of leading baselines for clarity; comprehensive comparisons against the fourteen methods, as well as additional 2K and 4K results, are provided in the Appendix. At 4K resolution, our method renders anatomically correct hands with exceptional detail as shown in the first row, while SANA produces a blurry image, HiFlow introduces diagonal artifacts, and I-MAX fails to generate the correct number of fingers. In the second row, UltraGen consistently preserves textural clarity, such as tree structures, surpassing competitors while maintaining global coherence. The third row highlights highly detailed eye and jewelries generation at 4K, where our method clearly captures sharpe details of eyelashes, and jewelries with fine structure. Training free methods like HiFlow [3] in the fourth row frequently shows diagonal artifacts, while another efficient attention method [28] failed to preserve global consistency as shown in the fifth columns. These examples collectively demonstrate UltraGen’s ability to synthesize images that combine fine-grained local details with coherent large-scale composition, validating its performance across diverse high-resolution scenarios.
@@ 4.2 Comparison with State-of-the-Art Methods
Table 1: Quantitative comparison at 4K × 4K resolution. The best result is highlighted in bold, while the second-best result is underlined.  UltraGen consistently perform competitive results even compared to training-based methods. We also include our work with stronger base model FLUX2 here (in Italic) for reference.
@@
SANA 0.4457 76.31 74.27 16.68 14.02 0.3197
Diffusion-4K 0.3012 121.85 120.59 14.39 10.77 0.2844
UltraPixel 0.4421 77.42 70.94 16.98 13.26 0.3251
URAE 0.4369 67.39 62.56 17.11 12.39 0.3204
FLUX+BSRGAN 0.3897 71.39 63.45 17.08 12.87 0.3210
UltraFlux 0.4371 75.89 65.12 17.01 13.24 0.3165
DemoFusion 0.4392 74.89 66.37 16.23 13.02 0.3187
DiffuseHigh 0.4221 81.54 73.35 16.15 13.08 0.3175
HiDiffusion 0.4021 172.46 217.49 9.43 8.12 0.3129
FreCas 0.2704 213.44 322.08 7.98 6.84 0.2805
I-MAX 0.4381 70.33 65.67 16.50 12.69 0.3211
HiFlow 0.4407 69.18 63.72 17.13 13.43 0.3113
CLEAR 0.3463 131.15 120.79 14.18 11.92 0.2913
CLEAR+SDEdit 0.4352 88.23 87.19 16.84 13.78 0.3221
UltraGen 0.4505 67.03 61.78 17.21 13.31 0.3231
UltraGen(FLUX2) 0.4594 61.38 59.82 18.02 14.93 0.3267
Quantitative Comparison. Table 1 presents a quantitative comparison of our method against the nine state-of-the-art methods at 4K resolution. The results show that our method achieves highly competitive performance across all metrics, even outperforming training-based methods that rely on extensive 4K-resolution training data. Table 2 demonstrates consistent performance on the GenEval [15] benchmark, confirming that UltraGen’s output quality is resolution-agnostic.
Efficiency and Scalability Analysis. The efficiency gains of our approach are summarized in Figure 4 and Figure 5. For a fair comparison, we focus on methods built upon the FLUX.1-dev backbone without architectural modifications. Under this setting, our method achieves a $1.7\times$ to $4.3\times$ inference speedup over its counterparts (with all methods exhibiting greater than $2.3\times$ speedup per denoising iteration). Although CLEAR [28] (augmented with SDEdit [32]) attains a comparable per-iteration speed, our method surpasses it in overall inference time by substantially reducing the number of required denoising steps. In terms of memory consumption, our method is highly efficient: excluding the base model’s footprint $\sim 32\,GB)$, it requires only 7.3GB7.3\,GB of additional VRAM, whereas a naive implementation with FlashAttention consumes 15.6GB15.6\,GB. This significant reduction in memory overhead enables inference on consumer-level GPUs (i.e., 40/48GB40/48\,GB) without module offloading, and facilitates more GPU-friendly fine-tuning.
Figure 5: Efficiency comparisons across resolutions (1K to 8K). We report per-iteration latency (Left), floating-point operations per attention layer (Middle), and total inference time (Right) for our method, naive matrix multiplication attention, Linear Attention, FlashAttention-2, and two linearlized attention configurations from CLEAR. At 8K resolution, our method achieves a $6.93\times$ per-iteration speedup, $130.4\times$ reduction in attention computation, and $10.77\times$ total inference speedup over FLUX with FlashAttention, while consistently outperforming CLEAR and Linear Attention across all metrics. The "X" mark means out-of-memory (OOM). Linear Attention reaches OOM at 8K resolution, while the naive one reaches OOM at 3K resolution.
@@
Figure 5 further illustrates the scalability of our approach. As the target resolution increases from 1K to 8K, our method achieves a $10.77\times$ reduction in latency and a $130.4\times$ reduction in floating-point computation within the attention layer compared to FlashAttention-2 [8]. Compared with another linearized attention mechanism [28], our approach remains consistently faster and more memory-efficient. We select two optimal configurations from their paper—CLEAR with a window radius of 32 and a radius of 16 with a downsample factor of 4—and as shown in the line chart, our method yields faster inference and lower computational cost in the attention layer as resolution scales, while also achieving superior generation quality as demonstrated in Table 1 and Figure 3. We attribute these efficiency gains to two key design choices: (1) our dense-block-style local window attention is more hardware-friendly, with window sizes that align well with GPU architectures; and (2) we fully leverage low-resolution guidance as an initialization signal, in contrast to CLEAR, which employs low-resolution guidance solely as a starting point for SDEdit due to its limited scalability across resolutions. For linear attention, while also having relatively low GFLOPs, it requires scaling the hidden dimension with sequence length to maintain performance [17, 33, 65], and it is not able to produce meaningful results via directly substituting the softmax based attentions in pretrained models [28]. In contrast, our method uses local window attention, which does not need to scale with resolution, enabling consistent memory and latency efficiency as resolution increases. Since linear attention requires the hidden dimension to scale with resolution, previous works would need to fine-tune on high-resolution data for each target resolution, whereas our framework achieves ultra-high-resolution synthesis using only 1K-resolution training data without any performance drop. Collectively, these results validate the scalability, efficiency, and robustness of our proposed framework.
@@ 5 Ablation Study
Window size and resolution ratio. We conduct ablation study on window size (ws) and ratio between high-res and low-res image ($\rho$). We experiment with window sizes from 256, 512, and 1024, and $\rho=2,4$. As shown in Table 3, the variations in FiD scores across these configurations were not statistically significant. To maximize computational efficiency without compromising performance, we selected the smallest window size (ws=256) and the largest ratio ($\rho$=4). It is worth noting that smaller window sizes or larger ratios are constrained by the fixed granularity of the kernel which depends on the hardware design of the GPU.
@@
Robustness on Different Models To demonstrate the robustness of UltraGen, we conducted additional experiments on FLUX.2-series. We provide both qualitative (Figure 3 and Figure 1) and quantitative (Figure 1). (Table 1 results for FLUX.2-Klein-9B, which is a time-distilled 4 step model. We therefore achieve sub 30-second generation for 4K images and sub 2-minute for 8K images, enabling maximum efficiency in ultra-high-resolution content creation.
Content in the Supplementary Materials. We will include (1) More 8K, 4K and 2K resolution galleries, and the PDF file for the Teasor figure. (2) a detailed quantitative 4K resolution comparison against all 14 baselines, (3) demo for our method’s compatability with different styled LoRA models (4) more ablation on the low-resolution initialization, and (5) discussion of why our method can generalize to higher-resolution than pretraining.
@@ 6 Conclusion
We present UltraGen, a novel framework enabling pre-trained diffusion transformers to generate ultra-high-resolution images without requiring additional high-resolution training data. UltraGen introduces a hierarchical attention mechanism that partitions high-resolution latents into fixed-size local windows while maintaining global coherence through low-resolution guidance with scaled positional anchors. The framework comprises three key components: (i) efficient local window attention that reduces quadratic complexity to near-linear scaling; (ii) global semantic preservation via low-resolution guidance latents with position scaling; and (iii) parameter-efficient joint denoising through LoRA-based adaptations trained solely on commodity resolutions. Extensive experiments demonstrate that UltraGen achieves superior visual quality at 4K resolution while delivering over $10\times$ speedup in $8K$ resolution and less memory usage compared to dense attention baselines, establishing a practical paradigm for scalable ultra-high-resolution text-to-image generation.
Supplementary Materials
Table 4: Ablation study on the number of skipped denoising steps at 4K resolution. PSNR is computed against the full-step (no skipping) output.
Skipped Steps 0 (Ref.) 4 8 10 14 18
PSNR (dB) $\uparrow$ 28.78 28.74 28.70 27.43 26.43
[…]
7 Ultra-High-Res Gallery (>4K>4K)
In Figure 7, we show one $8192\times 8192$ and one $4096\times 10240$ images, more demos are in the Zip file in the supplementary materials.
Figure 7: One $8192\times 8192$ and One $4096\times 10240$ images
8 4K and 2K Galleries
In Figure 1089 and Figure 12 we present more 4K and 2K resolution results. We also provide the PDF for the Teasor figure.
@@ 7 Ultra-High-Res Gallery (>4K>4K)
Figure 10: More 4K results
Figure 11: 4K results with styled LoRAs, zoom in to see the details.
@@
9 Complete Qualitative Comparison
In figure 13, we present the complete qualitative comparison of 4K resolution images with the quantitatively compared baselines in the main paper. We provide the PDF in the Zip file.
Figure 13: Complete comparison against all baselines on 4K resolution. Zoom in to view the details, our method produces the best results.
10 Results with Different Styled LoRAs
Our model is compatible with other styled LoRAs, Figure 11 present 4K results from 3 different LoRAs, specifically "XLabs-AI/flux-RealismLora",
"prithivMLmods/Flux-Product-Ad-Backdrop",
and "UmeAiRT/FLUX.1-dev-LoRA-Modern_Pixel_art".
11 Ablation Study on Step Skipping.
We conduct an ablation study to investigate the optimal number of skippable denoising steps, seeking a favorable trade-off between reconstruction quality and inference speed. Specifically, we evaluate configurations in which 18, 14, 10, 8, and 4 steps are skipped during the denoising process at 4K resolution. Peak Signal-to-Noise Ratio (PSNR) is adopted as the quantitative metric, with the full-step (no skipping) output serving as the ground-truth reference. The results are summarized in Table 4. In practice, we also find that by skipping 10 steps, we can have the most natural results as well as comparable sharpness and if too many steps are skipped like >14>14 steps, the denoising will lean towards oversmooth results as the low-resolution images.
12 Discussion on Advantages of Window Permuted Local Attention
12.1 Quality Preservation Across Resolutions
A critical advantage of our approach is that representational quality remains constant regardless of output resolution. Since every local window operates on the same relative position range $\mathcal{R}_{\text{local}}=\{(\Delta i,\Delta j):|\Delta i|,|\Delta j|\leq w-1\}$, and this range falls entirely within the pretrained distribution $\mathcal{D}_{\text{train}}$, each window achieves identical quality. For an image of size $H\times W$ with local windows of size $w\times w$, and $s$ is the VAE downsampling factor The per-unit quality across the entire image is: